OpenStudy (anonymous):

what is the closest benchmark percent to 33 over 60?

OpenStudy (astrophysics):

33/60 = 0.55*100 = 55%

OpenStudy (anonymous):

the answers are 75% 50% 25% and 0%

OpenStudy (astrophysics):

So what number will be closest? :P

OpenStudy (anonymous):

50%

OpenStudy (astrophysics):

Yes

OpenStudy (anonymous):

thanks